Circular No. 5764

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence; and

R. M. Wagner, Ohio State University, report: “We confirm that a rapid decrease in brightness of the optical counterpart of GRO J0422+32 took place during the first half of April. A spectrum taken with the 1.8-m Perkins telescope at Lowell Observatory on Apr. 7.1 UT showed Hα, Hβ, and very weak He II 468.6-nm emission, with the object being at mag ∼ 15-16. CCD photometry obtained with the 1.2-m Haute-Provence telescope on Apr. 8.8 and 11.8 UT yielded V = 16.2 and 16.6, respectively. Our observations, combined with the estimate of Filippenko and Matheson (IAUC 5761), indicate a final decline rate of 0.3-0.4 mag/day. The total duration of the optical outburst of this source (∼ 250 days since x-ray turn-on) is remarkably close to that observed in V616 Mon (A0620–00) in 1975-76.”

GRB 930309 A. J. Beasley, K. S. Dwarakanath, and M. P. Rupen, National Ra-

dio Astronomy Observatory, report: “VLA snapshot observations on Apr. 9.93 (1425 MHz) and 14.93 UT (330 and 1425 MHz) of the radio source previously associated with GRB 930309 (IAUC 5749, 5750) indicate an extended source consisting of two lobe-like components. The components are located at α = 21^{h}31^{m}16^{s}.3, δ = +54^{o}26′55′′.4 and 21^{h}32^{m}14^{s}.6, +54^{o}27′11′′.9 (equinox 1950.0), with integrated 1425-MHz flux densities of 9.7 ± 0.2 and 6.6 ± 0.2 mJy, respectively (1σ errors). There is no evidence for significant variation in either component between the 1425-MHz observations. At 330 MHz, we detect an extended source with an integrated flux density of 154.9 ± 6 mJy, implying a combined spectral index α ∼ –1.5 (ν^{α}). This index is also consistent with the 49-cm flux density and 6-cm upper limit reported on IAUC 5749. The source appears to be an extragalactic background source, presumably unrelated to GRB 930309 (cf. IAUC 5763).”

S APODIS P. M. Kilmartin, A. C. Gilmore, and P. L. Cottrell, Mt. John University

Observatory, provide the following photoelectric measures of this fading R CrB-type star (all measures ± 0.02 mag): Mar. 28.48 UT, V = 9.87, B–V = +1.15; Apr. 1.39, 10.03, +1.19; 5.40, 10.22, +1.22; 6.39, 10.28, +1.19; 11.46, 10.61, +1.22; 14.44, 10.88, +1.24.

1993 April 16 Daniel W. E. Green